Shaypop:

AO= 5x+2 BO= x-1 CO=3x-7 DO= 3x+8 Quadrilateral ABCD is a parallelogram if it's diagonal bisect each other. Prove that quadrilateral parallelogram by finding the value of x.

Shaypop:

jhonyy9:

do you can post an image about this parallelogram ?

Shaypop:

jhonyy9:

do you know the properties of a parallelogram ? how are the opposite angles and sides ?

Shaypop:

Yes but I don't study

Shaypop:

Come on

Shaypop:

@AZ

Shaypop:

HELP!!!

Shaypop:

I'm finna cry

jhonyy9:

the diagonales bisect each other - what mean this ? how you can write this using the given details ?

Shaypop:

Add up to 180

jhonyy9:

no look please on this image how are OB and OC or OA and OD ?

jhonyy9:

signed on this image

Shaypop:

Divide by 2

jhonyy9:

no yes the diagonales bisect each other

jhonyy9:

from this result OA = OD and OB = OC

Shaypop:

90 divided by 2

jhonyy9:

rewrite these equalities using these given expressions

Shaypop:

They equal to each other

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shaypop 90 divided by 2 \(\color{#0cbb34}{\text{End of Quote}}\) you dont work now with angles

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 from this result OA = OD and OB = OC \(\color{#0cbb34}{\text{End of Quote}}\) using these given expressions rewrite these equalities

Shaypop:

5x+2=3x+2

jhonyy9:

OA=OD 5x +2 = 3x +8

Shaypop:

3

jhonyy9:

what is this 3 ?

Shaypop:

The answer

jhonyy9:

do you think x = 3 ?

Shaypop:

No

jhonyy9:

why no ? you need find the x value and prove that diagonales bisect each other

Shaypop:

It is b

jhonyy9:

what is this b ? i dont see there given any choices for x value

jhonyy9:

x = ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 OA=OD 5x +2 = 3x +8 \(\color{#0cbb34}{\text{End of Quote}}\) use this to calcule x = ?

Shaypop:

It's 3

jhonyy9:

so x = 3 then now can you prove that the diagonales bisect each other ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 OA=OD 5x +2 = 3x +8 \(\color{#0cbb34}{\text{End of Quote}}\) use this

jhonyy9:

substitute in place of x the 3

Shaypop:

I did

jhonyy9:

and what you get ?

Shaypop:

2,2 and 17,17

jhonyy9:

so what mean this ? what result from these ?

Shaypop:

I plug it in

jhonyy9:

written in this way 2 = 2 17 = 17

Shaypop:

Ok

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 written in this way 2 = 2 17 = 17 \(\color{#0cbb34}{\text{End of Quote}}\) from these result that the diagonales bisect each other so result that ABCD is a parallelogram ok ?

jhonyy9:

questions ?